OpenStudy (anonymous):

A box contains 4 black checkers and 8 red checkers. A checker is drawn and replaced. Then a second checker is drawn. Find the probability that both checkers are red.

OpenStudy (anonymous):

multiply the probability that the first one is red, times the probability that the second one is red

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{8}{12}\times \frac{8}{12}\] or maybe \[\frac{2}{3}\times \frac{2}{3}\]

OpenStudy (anonymous):

for the first red the prob is 8/12 and for the second too it is 8/12 (note replacement done) hence th reqd probability is 8/12 *8/12=4/9

OpenStudy (anonymous):

ya it's 4/9 that's what I got thanks guys